# Approximation of magnetic Schrödinger operators with $\delta$-interactions supported on networks

### v1

@@ 2.3. Rigorous definition of $H_{A,Q,\alpha}$
The goal in this subsection is to study the form $\mathfrak{h}_{A,Q,\alpha}$ from (1.7), which then leads to a rigorous definition of the formal expression in (1.1).
@@
$\|f\|_{\mathcal{H}_{A,Q}}^{2}\leq C_{1}(\textup{Re}\,\mathfrak{h}_{A,Q,\alpha}+C_{2})[f]$ (2.11)
@@
By assumption, for each $k\in\{1,\dots,N\}$ there exist functions $\alpha^{(k)}_{p}\in L^{p}(\Sigma^{(k)})$ and $\alpha^{(k)}_{\infty}\in L^{\infty}(\Sigma^{(k)})$ such that $\alpha^{(k)}=\alpha^{(k)}_{p}+\alpha^{(k)}_{\infty}$. Set $\alpha_{p}:=(\alpha^{(k)}_{p})_{k=1}^{N}\in L^{p}(\Sigma)$ and $\alpha_{\infty}:=(\alpha^{(k)}_{\infty})_{k=1}^{N}\in L^{\infty}(\Sigma)$, so that $\alpha=\alpha_{p}+\alpha_{\infty}$. Define for $*\in\{p,\infty\}$ the form
@@
Let $A\in L^{2}_{\textup{loc}}(\mathbb{R}^{n};\mathbb{R}^{n})$ and $Q=Q_{1}+Q_{2}$ with $0\leq Q_{1}\in L^{1}_{\textup{loc}}(\mathbb{R}^{n})$ and $Q_{2}\in L^{p^{\prime}}(\mathbb{R}^{n})$, where $p^{\prime}\geq\frac{n}{2}$ for $n>2$ and $p^{\prime}>1$ for $n=2$. Moreover, let $\Sigma$ be an admissible $C^{2}$-hypersurface with unit normal vector field $\nu$, $q=2\frac{n-1}{n-2+\mu}$ for $\mu\in(0,1]$, and $\mu^{\prime}\in(0,\mu]\cap(0,1)$. Then, there exists $\beta_{0}>0$ and a constant $C_{1}>0$ such that for all $r\in(-\beta_{0},\beta_{0})$ the estimate
@@
is true. In particular, there exists a constant $C_{2}>0$ such that for all $r\in(-\beta_{0},\beta_{0})$
@@
Step 1. Let $\widetilde{\nu}$ be a $C^{1}_{b}$ extension of $\nu$ to $\mathbb{R}^{n}$. To show that such an extension exists, let $\Omega$ be as in Definition 2.5 and $\widehat{\nu}$ be the normal vector field at $\partial\Omega$; then, as $\Sigma\subset\partial\Omega$, the function $\widehat{\nu}$ is an extension of $\nu$ to $\partial\Omega$. Moreover, choose $\phi\in C^{1}_{b}(\mathbb{R};\mathbb{R})$ with $\phi(0)=1$ and $\textup{supp}\,\phi\subset(-\beta^{\Omega},\beta^{\Omega})$, where $\beta^{\Omega}$ is the number from Proposition 2.6 for the admissible hypersurface $\partial\Omega$. Then one possible choice of $\widetilde{\nu}$ is given by
@@
First, we show (3.3) for $\delta=1$. With the main theorem of calculus, one finds for $f\in C_{0}^{\infty}(\mathbb{R}^{n})$ that
@@
As in the proof of [5, Theorem 4.5], see equation (B.9) there, one can show that for $q=2$ the estimate in (3.1) is true for $\mu^{\prime}=1$. This can be used to show that the difference in (1.12) can be estimated by $C\varepsilon^{1/2}$, if $V_{\varepsilon}\in L^{\infty}(\mathbb{R}^{n})$. Since a main focus is on more singular potentials $V_{\varepsilon}$, this small possible improvement is not further investigated in this paper.
@@
$\begin{split}\mathfrak{a}[f]&=\sum_{k=1}^{N}\left[\int_{\Omega_{\varepsilon}^{(k)}}V_{\varepsilon}^{(k)}(x)|f(x)|^{2}\text{d}x-\int_{\Sigma^{(k)}}\alpha(x_{\Sigma^{(k)}})|f(x_{\Sigma^{(k)}})|^{2}\text{d}\sigma(x_{\Sigma^{(k)}})\right]\\ &=\sum_{j=1}^{N}\bigg{[}\int_{\Sigma^{(k)}}\int_{-\varepsilon}^{\varepsilon}\frac{\beta}{\varepsilon}V^{(k)}\bigg{(}\iota^{(k)}\bigg{(}x_{\Sigma^{(k)}},\frac{\beta}{\varepsilon}t\bigg{)}\bigg{)}|f(\iota^{(k)}(x_{\Sigma^{(k)}},t))|^{2}\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\cdot\det(1-tW^{(k)}({x_{\Sigma^{(k)}}}))\text{d}t\text{d}\sigma(x_{\Sigma^{(k)}})\\ &\qquad\qquad-\int_{\Sigma^{(k)}}\int_{-\beta}^{\beta}V^{(k)}(\iota^{(k)}(x_{\Sigma^{(k)}},r))\text{d}r|f(\iota^{(k)}(x_{\Sigma^{(k)}},0))|^{2}\text{d}\sigma(x_{\Sigma^{(k)}})\bigg{]}\\ &=\sum_{j=1}^{N}\int_{\Sigma^{(k)}}\int_{-\beta}^{\beta}V^{(k)}\big{(}\iota^{(k)}(x_{\Sigma^{(k)}},r)\big{)}\\ &\qquad\qquad\qquad\cdot\bigg{[}\bigg{|}f\bigg{(}\iota^{(k)}\bigg{(}x_{\Sigma^{(k)}},\frac{\varepsilon}{\beta}r\bigg{)}\bigg{)}\bigg{|}^{2}\det\bigg{(}1-\frac{\varepsilon}{\beta}rW^{(k)}({x_{\Sigma^{(k)}}})\bigg{)}\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\big{|}f(\iota^{(k)}(x_{\Sigma^{(k)}},0))\big{|}^{2}\bigg{]}\text{d}t\text{d}\sigma(x_{\Sigma^{(k)}}),\\ \end{split}$
where the substitution $r=\frac{\beta}{\varepsilon}t$ was done in the last step. Since by assumption $V^{(k)}\in L^{p}(\mathbb{R}^{n})+L^{\infty}(\mathbb{R}^{n})$, there exist $V^{(k)}_{p}\in L^{p}(\mathbb{R}^{n})$ and $V_{\infty}^{(k)}\in L^{\infty}(\mathbb{R}^{n})$ such that $V^{(k)}=V_{p}^{(k)}+V_{\infty}^{(k)}$. As $\textup{supp}\,V^{(k)}\subset\Omega_{\beta}^{(k)}$, it is no restriction to assume for $*\in\{p,\infty\}$ that also $\textup{supp}\,V_{*}^{(k)}\subset\Omega_{\beta}^{(k)}$. Define for $*\in\{p,\infty\}$
@@ 4. Consequences for the spectrum of $H_{A,Q,V_{\varepsilon}}$
Since convergence in the norm resolvent sense implies the convergence of the spectra of the associated operators, Theorem 1.1 allows to transfer results on the spectrum of $H_{A,Q,\alpha}$ to the spectrum of $H_{A,Q,V_{\varepsilon}}$. If $\Sigma$ is one $C^{2}$-hypersurface, some interesting consequences are elaborated in [4, Section 3.3]; in particular, in [4, Proposition 3.5] results on the existence of eigenvalues of $H_{0,0,\alpha}$ (for $A=Q=0$) in various situations are transferred to $H_{0,0,V_{\varepsilon}}$. In this section, some further consequences are discussed, which can not be obtained from the results in [4] due to the different assumptions there. Here, the focus lies in self-adjoint examples; however, since convergence in the norm resolvent sense also implies the convergence of the spectrum in the non self-adjoint setting, see for instance [27, Section IV.3.5], also results on the spectrum of $H_{A,Q,\alpha}$ for non-symmetric $\alpha$ from, e.g., [10, 24, 28] could be transferred.
@@
Let $A$ and $Q=Q_{1}+Q_{2}$ be as in (1.3) and (1.4), respectively, $\Sigma$ be as in (1.2), and $V^{(k)}\in L^{p}(\mathbb{R}^{n})+L^{\infty}(\mathbb{R}^{n})$ with $p=\frac{n-1}{1-\gamma}$ and $\gamma\in(0,1)$ be supported in $\Omega_{\beta}^{(k)}$. Assume additionally that $Q$ and $V^{(k)}$ are real-valued. Let for $\varepsilon_{0}$ as in Theorem 1.1 the potential $V_{\varepsilon}$ be defined by (1.9) and $H_{A,Q,V_{\varepsilon}}$ be the self-adjoint operator associated with the form $\mathfrak{h}_{A,Q,V_{\varepsilon}}$ in (1.10). Eventually, let $\alpha$ be as in (1.11) and $H_{A,Q,\alpha}$ be the self-adjoint operator associated with the form $\mathfrak{h}_{A,Q,\alpha}$ in (1.7). Then, the following is true:
@@
    $\lambda\in\sigma_{\textup{disc}}(H_{A,Q,\alpha})$ if and only of there exists a family of numbers $(\lambda_{\varepsilon})_{\varepsilon\in(0,\varepsilon_{0})}$ with $\lambda_{\varepsilon}\in\sigma_{\textup{disc}}(H_{A,Q,V_{\varepsilon}})$ such that $\lambda_{\varepsilon}\rightarrow\lambda$.
@@
    $\lambda\in\sigma_{\textup{ess}}(H_{A,Q,\alpha})$ if and only of there exists a family of numbers $(\lambda_{\varepsilon})_{\varepsilon\in(0,\varepsilon_{0})}$ with $\lambda_{\varepsilon}\in\sigma_{\textup{ess}}(H_{A,Q,V_{\varepsilon}})$ such that $\lambda_{\varepsilon}\rightarrow\lambda$.
@@
Note that this $\Sigma$ falls into the class of networks considered in this paper, as it is the union of two rays $\Sigma^{(1)}$ and $\Sigma^{(2)}$ with joint endpoint at the origin.
@@
$F_{\phi,\alpha}(x,y):=1+\frac{x^{4}}{4}-x^{2}\Theta_{\delta,\phi}+\frac{\alpha}{\sqrt{\pi}}xe^{-y^{2}\tan^{2}(\phi/2)}(1+\textup{erf}(y)),$
where $\Theta_{\delta,\phi}=\inf\sigma_{\textup{ess}}(H_{A,0,\alpha})>-\infty$ and erf is the error function. If $\varepsilon>0$ is sufficiently small and $\inf_{x,y\in(0,\infty)}F_{\phi,\alpha}(x,y)<0$, then
@@
First, taking the given assumptions into account, there exists a point $\lambda_{0}\in\sigma_{\textup{disc}}(H_{A,0,\alpha})\cap(-\infty,\Theta_{\delta,\phi})$, see [20, Theorem 3 and Corollary 2] (note that $\beta$ in [20] is equal to $-\alpha$). Since by Lemma 4.1 one has $\inf\sigma_{\textup{ess}}(H_{A,0,V_{\varepsilon}})\rightarrow\Theta_{\delta,\phi}$ and there exists $\lambda_{\varepsilon}\in\sigma_{\textup{disc}}(H_{A,0,V_{\varepsilon}})$ such that $\lambda_{\varepsilon}\rightarrow\lambda_{0}$, as $\varepsilon\searrow 0$, one concludes that $\lambda_{\varepsilon}<\inf\sigma_{\textup{ess}}(H_{A,0,V_{\varepsilon}})$. ∎
@@
To formulate the mentioned optimization result for Schrödinger operators with potentials supported in a tubular neighborhood of a star graph, let $N\in\mathbb{N}$, $N>2$, be fixed and $\phi_{1},\dots\phi_{N}\in(0,2\pi)$ such that $\phi_{1}+\dots+\phi_{N}=2\pi$. Let $\Sigma\subset\mathbb{R}^{2}$ be a star graph, where each of the edges has length $L\in(0,\infty)$, i.e.
$\Sigma=\{(r\cos\theta,r\sin\theta):r\in(0,L),\theta\in\{\phi_{1},\dots,\phi_{N}\}\}.$
Moreover, denote by $\Gamma$ the symmetric star graph with $N$ edges of length $L$, i.e. $\Gamma$ is as $\Sigma$ with $\phi_{1}=\dots=\phi_{N}=\frac{2\pi}{N}$. Note that both graphs $\Sigma$ and $\Gamma$ are networks as in (1.2), as they are unions of $N$ line segments $\Sigma^{(1)},\dots,\Sigma^{(N)}$ and $\Gamma^{(1)},\dots,\Gamma^{(N)}$, respectively. Let $V_{\varepsilon}$ be as described after Lemma 4.1 and let $W_{\varepsilon}$ be constructed in the same way, but with the network $\Sigma$ being replaced by $\Gamma$ with the same $\alpha$ and the same $\beta$ as in the definition of $V_{\varepsilon}$, and denote the associated Schrödinger operator by $H_{0,0,W_{\varepsilon}}$. As $\Sigma$ and $\Gamma$ are compact, $V_{\varepsilon}$ and $W_{\varepsilon}$ are bounded and compactly supported and thus $\sigma_{\textup{ess}}(H_{0,0,V_{\varepsilon}})=\sigma_{\textup{ess}}(H_{0,0,W_{\varepsilon}})=[0,\infty)$. One has the following optimization result for the lowest eigenvalue of $H_{0,0,V_{\varepsilon}}$.
@@
First, by [19, Theorem 1.1] the Schrödinger operators $H_{0,0,\alpha}(\Sigma)$ and $H_{0,0,\alpha}(\Gamma)$ with $\delta$-potentials supported on $\Sigma$ and $\Gamma$, respectively, have negative discrete eigenvalues and for their smallest eigenvalues $\lambda_{0}(\Sigma)$ and $\lambda_{0}(\Gamma)$ the inequality
@@
Note that this curve with a cusp at the origin can be realized by gluing together two admissible curves $\Sigma^{(1)}$ and $\Sigma^{(2)}$ at the origin and hence, it falls into the class of networks considered in the present paper. To formulate the next result, one has to introduce in $L^{2}(0,\infty)$ the differential operator

### v2

@@ 2.3. Rigorous definition of $H_{A,Q,\alpha}$
The goal in this subsection is to study the form $\mathfrak{h}_{A,Q,\alpha}$ from (1.7), which then leads to a rigorous definition of the formal expression in (1.1). The proof of the following proposition follows similar ideas as in the construction of non-magnetic Schrödinger operators with $\delta$-potentials, see, for instance, [11, Proposition 2.1], [14, Section 4] or [4, Lemma 2.7].
@@
$\|f\|_{\mathcal{H}_{A,Q}}^{2}\leq C_{1}\big(\textup{Re}\,\mathfrak{h}_{A,Q,\alpha}[f]+C_{2}\|f\|_{L^{2}(\mathbb{R}^{n})}^{2}\big)$ (2.11)
@@
By assumption, for each $k\in\{1,\dots,N\}$ there exist functions $\alpha^{(k)}_{p}\in L^{p}(\Sigma^{(k)})$ and $\alpha^{(k)}_{\infty}\in L^{\infty}(\Sigma^{(k)})$ such that $\alpha^{(k)}=\alpha^{(k)}_{p}+\alpha^{(k)}_{\infty}$. Set $\alpha_{p}:=(\alpha^{(k)}_{p})_{k=1}^{N}\in L^{p}(\Sigma)$ and $\alpha_{\infty}:=(\alpha^{(k)}_{\infty})_{k=1}^{N}\in L^{\infty}(\Sigma)$, so that $\alpha=\alpha_{p}+\alpha_{\infty}$. Define for $\in\{p,\infty\}$ the form
@@
Let $A\in L^{2}_{\textup{loc}}(\mathbb{R}^{n};\mathbb{R}^{n})$ and $Q=Q_{1}+Q_{2}$ with $0\leq Q_{1}\in L^{1}_{\textup{loc}}(\mathbb{R}^{n})$ and $Q_{2}\in L^{p^{\prime}}(\mathbb{R}^{n})$, where $p^{\prime}\geq\frac{n}{2}$ for $n>2$ and $p^{\prime}>1$ for $n=2$. Moreover, let $\Sigma$ be an admissible $C^{2}$-hypersurface with unit normal vector field $\nu$, $q=2\frac{n-1}{n-2+\mu}$ for $\mu\in(0,1]$, and $\mu^{\prime}\in(0,\mu]\cap(0,1)$. Then, there exists $\beta_{0}>0$ and a constant $C_{1}>0$ such that for all $r\in(-\beta_{0},\beta_{0})$ and all $f\in\mathcal{H}_{A,Q}$ the estimate
@@
is true. In particular, there exists a constant $C_{2}>0$ such that for all $r\in(-\beta_{0},\beta_{0})$ and all $f\in\mathcal{H}_{A,Q}$
@@
Step 1. Let $\widetilde{\nu}$ be a $C^{1}_{b}$-extension of $\nu$ to $\mathbb{R}^{n}$. To show that such an extension exists, let $\Omega$ be as in Definition 2.5 and $\widehat{\nu}$ be the normal vector field at $\partial\Omega$; then, as $\Sigma\subset\partial\Omega$, the function $\widehat{\nu}$ is an extension of $\nu$ to $\partial\Omega$. Moreover, choose a real-valued function $\phi\in C^{1}_{b}(\mathbb{R})$ with $\phi(0)=1$ and $\textup{supp}\,\phi\subset(-\beta^{\Omega},\beta^{\Omega})$, where $\beta^{\Omega}$ is the number $\beta$ from Proposition 2.7 for the admissible hypersurface $\partial\Omega$. Then one possible choice of $\widetilde{\nu}$ is given by
@@
First, (3.3) is shown for $\delta=1$. With the main theorem of calculus, one finds for $f\in C_{0}^{\infty}(\mathbb{R}^{n})$ that
@@
As in the proof of [5, Theorem 4.5], see equation (B.9) there, one can show that for $q=2$ the estimate in (3.1) is true for $\mu^{\prime}=1$. This can be used to show that the difference in (1.12) can be estimated by $C\varepsilon^{1/2}$, if $V_{\varepsilon}\in L^{\infty}(\mathbb{R}^{n})$. Since a main focus in this paper is on more singular potentials $V_{\varepsilon}$, this small possible improvement is not further investigated here.
@@
[…]
where the substitution $r=\frac{\beta}{\varepsilon}t$ was done in the last step. Since by assumption $V^{(k)}\in L^{p}(\mathbb{R}^{n})+L^{\infty}(\mathbb{R}^{n})$, there exist $V^{(k)}_{p}\in L^{p}(\mathbb{R}^{n})$ and $V_{\infty}^{(k)}\in L^{\infty}(\mathbb{R}^{n})$ such that $V^{(k)}=V_{p}^{(k)}+V_{\infty}^{(k)}$. As $\textup{supp}\,V^{(k)}\subset\Omega_{\beta}^{(k)}$, it is no restriction to assume for $\in\{p,\infty\}$ that also $\textup{supp}\,V_{*}^{(k)}\subset\Omega_{\beta}^{(k)}$. Define for $\in\{p,\infty\}$
@@ 4. Consequences for the spectrum of $H_{A,Q,V_{\varepsilon}}$
Since convergence in the norm resolvent sense implies the convergence of the spectra of the associated operators, Theorem 1.1 allows to transfer results on the spectrum of $H_{A,Q,\alpha}$ to the spectrum of $H_{A,Q,V_{\varepsilon}}$. If $\Sigma$ is one $C^{2}$-hypersurface, some interesting consequences are elaborated in [4, Section 3.3]; in particular, in [4, Proposition 3.5] results on the existence of eigenvalues of $H_{0,0,\alpha}$ (for $A=Q=0$) in various situations are transferred to $H_{0,0,V_{\varepsilon}}$. In this section, some further consequences are discussed, which can not be obtained from the results in [4] due to the different assumptions there. Here, the focus lies in self-adjoint examples; however, since convergence in the norm resolvent sense also implies the convergence of the spectrum in the non self-adjoint setting, see for instance [27, Section IV.3.5], also results on the spectrum of $H_{A,Q,\alpha}$ for non-real $\alpha$ from, e.g., [10, 24, 28] could be transferred.
@@
Let $A$ and $Q=Q_{1}+Q_{2}$ be as in (1.3) and (1.4), respectively, $\Sigma$ be as in (1.2), and $V^{(k)}\in L^{p}(\mathbb{R}^{n})+L^{\infty}(\mathbb{R}^{n})$ with $p=\frac{n-1}{1-\gamma}$ and $\gamma\in(0,1)$ be supported in $\Omega_{\beta}^{(k)}$. Assume additionally that $Q$ and $V^{(k)}$ are real-valued. Let for $\varepsilon\in(0,\varepsilon_{0})$ with $\varepsilon_{0}$ as in Theorem 1.1 the potential $V_{\varepsilon}$ be defined by (1.9) and $H_{A,Q,V_{\varepsilon}}$ be the self-adjoint operator associated with the form $\mathfrak{h}_{A,Q,V_{\varepsilon}}$ in (1.10). Eventually, let $\alpha$ be as in (1.11) and $H_{A,Q,\alpha}$ be the self-adjoint operator associated with the form $\mathfrak{h}_{A,Q,\alpha}$ in (1.7). Then, the following is true:
@@
    $\lambda\in\sigma_{\textup{disc}}(H_{A,Q,\alpha})$ if and only of there exists a family of numbers $(\lambda_{\varepsilon})_{\varepsilon\in(0,\varepsilon_{0})}$ with $\lambda_{\varepsilon}\in\sigma_{\textup{disc}}(H_{A,Q,V_{\varepsilon}})$ such that $\lambda_{\varepsilon}\rightarrow\lambda$ for $\varepsilon\searrow 0$.
@@
    $\lambda\in\sigma_{\textup{ess}}(H_{A,Q,\alpha})$ if and only of there exists a family of numbers $(\lambda_{\varepsilon})_{\varepsilon\in(0,\varepsilon_{0})}$ with $\lambda_{\varepsilon}\in\sigma_{\textup{ess}}(H_{A,Q,V_{\varepsilon}})$ such that $\lambda_{\varepsilon}\rightarrow\lambda$ for $\varepsilon\searrow 0$.
@@
Note that this $\Sigma$ falls into the class of networks considered in this paper, as it is the union of two rays $\Sigma^{(1)}$ and $\Sigma^{(2)}$ with joint endpoint at the origin. The associated potential $V_{\varepsilon}$ is given by
$V_{\varepsilon}=\begin{cases}2\alpha\varepsilon^{-1}&\text{ in }\Omega_{\varepsilon}^{(1)}\cap\Omega_{\varepsilon}^{(2)}\\ \alpha\varepsilon^{-1}&\text{ in }(\Omega_{\varepsilon}^{(1)}\cup\Omega_{\varepsilon}^{(2)})\setminus(\Omega_{\varepsilon}^{(1)}\cap\Omega_{\varepsilon}^{(2)}),\\ 0&\text{ else;}\end{cases}$
note that there exists a constant $c=c(\phi)$ such that $\Omega_{\varepsilon}^{(1)}\cap\Omega_{\varepsilon}^{(2)}\subset B(0,c\varepsilon)$.
@@
$F_{\phi,\alpha}(x,y):=1+\frac{x^{4}}{4}-x^{2}\Theta_{\delta,\alpha}+\frac{\alpha}{\sqrt{\pi}}xe^{-y^{2}\tan^{2}(\phi/2)}(1+\textup{erf}(y)),$
where $\Theta_{\delta,\alpha}=\inf\sigma_{\textup{ess}}(H_{A,0,\alpha})>$-$\infty$ and erf is the error function. If $\varepsilon>0$ is sufficiently small and $\inf_{x,y\in(0,\infty)}F_{\phi,\alpha}(x,y)<0$, then
@@
First, taking the given assumptions into account, there exists a point $\lambda_{0}\in\sigma_{\textup{disc}}(H_{A,0,\alpha})\cap(-\infty,\Theta_{\delta,\alpha})$, see [21, Theorem 3 and Corollary 2] (note that $\beta$ in [21] is equal to $\alpha$). Since by Lemma 4.1 one has $\inf\sigma_{\textup{ess}}(H_{A,0,V_{\varepsilon}})\rightarrow\Theta_{\delta,\alpha}$, as $\varepsilon\searrow 0$, and there exists $\lambda_{\varepsilon}\in\sigma_{\textup{disc}}(H_{A,0,V_{\varepsilon}})$ such that $\lambda_{\varepsilon}\rightarrow\lambda_{0}$, as $\varepsilon\searrow 0$, one concludes that $\lambda_{\varepsilon}<\inf\sigma_{\textup{ess}}(H_{A,0,V_{\varepsilon}})$. ∎
@@
To formulate the mentioned optimization result for Schrödinger operators with potentials supported in a tubular neighborhood of a star graph, let $N\in\mathbb{N}$, $N>2$, be fixed and $\phi_{1},\dots\phi_{N}\in(0,2\pi)$ such that $\phi_{1}+\dots+\phi_{N}=2\pi$, and set $\widetilde{\phi}_{k}:=\sum_{j=1}^{k}\phi_{j}$, $k\in\{1,\dots,N\}$. Let $\Sigma\subset\mathbb{R}^{2}$ be a star graph, where each of the edges has length $L\in(0,\infty)$, i.e.
$\Sigma=\{(r\cos\theta,r\sin\theta):r\in(0,L),\theta\in\{\widetilde{\phi}_{1},\dots,\widetilde{\phi}_{N}\}\}.$
Moreover, denote by $\Gamma$ the symmetric star graph with $N$ edges of length $L$, i.e. $\Gamma$ is as $\Sigma$ with $\phi_{1}=\dots=\phi_{N}=\frac{2\pi}{N}$. Note that both graphs $\Sigma$ and $\Gamma$ are networks as in (1.2), as they are unions of $N$ line segments $\Sigma^{(1)},\dots,\Sigma^{(N)}$ and $\Gamma^{(1)},\dots,\Gamma^{(N)}$, respectively. Let $V_{\varepsilon}$ be as described after Lemma 4.1 and let $W_{\varepsilon}$ be constructed in the same way, but with the network $\Sigma$ being replaced by $\Gamma$ with the same $\alpha$ and the same $\beta$ as in the definition of $V_{\varepsilon}$, and denote the associated Schrödinger operator by $H_{0,0,W_{\varepsilon}}$. Note that, as in the previous result, $\max|V_{\varepsilon}|\leq N|\alpha|\varepsilon^{-1}$ and, with a constant $c=c(\phi_{1},\dots,\phi_{N})$, the area in which $|V_{\varepsilon}|\geq|\alpha|\varepsilon^{-1}$, is contained in $B(0,c\varepsilon)$; a similar statement holds for $W_{\varepsilon}$. As $\Sigma$ and $\Gamma$ are compact, $V_{\varepsilon}$ and $W_{\varepsilon}$ are bounded and compactly supported and thus $\sigma_{\textup{ess}}(H_{0,0,V_{\varepsilon}})=\sigma_{\textup{ess}}(H_{0,0,W_{\varepsilon}})=[0,\infty)$. One has the following optimization result for the lowest eigenvalue of $H_{0,0,V_{\varepsilon}}$.
@@
First, by [20, Theorem 1.1] the Schrödinger operators $H_{0,0,\alpha}(\Sigma)$ and $H_{0,0,\alpha}(\Gamma)$ with $\delta$-potentials supported on $\Sigma$ and $\Gamma$, respectively, have negative discrete eigenvalues and for their smallest eigenvalues $\lambda_{0}(\Sigma)$ and $\lambda_{0}(\Gamma)$ the inequality
@@
Note that this curve with a cusp at the origin can be realized by gluing together two admissible curves $\Sigma^{(1)}$ and $\Sigma^{(2)}$ at the origin and hence, it falls into the class of networks considered in the present paper. In this case, the associated potential $V_{\varepsilon}$ is given by
$V_{\varepsilon}=\begin{cases}2\alpha\varepsilon^{-1}&\text{ in }\Omega_{\varepsilon}^{(1)}\cap\Omega_{\varepsilon}^{(2)}\\ \alpha\varepsilon^{-1}&\text{ in }(\Omega_{\varepsilon}^{(1)}\cup\Omega_{\varepsilon}^{(2)})\setminus(\Omega_{\varepsilon}^{(1)}\cap\Omega_{\varepsilon}^{(2)}),\\ 0&\text{ else;}\end{cases}$
note that there exists a constant $c=c(\varepsilon)\in(0,1)$, $c(\varepsilon)\rightarrow 1$ for $\varepsilon\searrow 0$, such that $\Omega_{\varepsilon}^{(1)}\cap\Omega_{\varepsilon}^{(2)}$ is contained in the $\varepsilon$-neighborhood of $|y|=x^{d}$ for $x\in(0,c\varepsilon^{1/d})$; this is a larger set than in the previous two examples. To formulate the next result, one has to introduce in $L^{2}(0,\infty)$ the differential operator
@@
Acknowledgement.
The author thanks the anonymous reviewers for constructive feedback that helped to improve the paper as well as Jussi Behrndt and Petr Siegl for fruitful discussions.
Funding statement.
This research was funded in part by the Austrian Science Fund (FWF) 10.55776/P 33568-N. For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
Data availability statement
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Competing Interests
The author has no competing interests to declare that are relevant to the content of this article.